Find the current through a loop needed to create a maximum torque of 18 N⋅m. The loop has 132 square turns that are 18.7 cm on a side and is in a uniform 0.242-T magnetic field.


answer to 2 decimal places


1
Expert's answer
2021-05-27T12:29:42-0400

Gives

Torques"\tau=18Nm"

Side square (l)=18.7cm

Magnetic field (B)=0.242T

Magnetic moment


"M=NIA"

"\tau=M\times B"

"\tau=NIABsin\theta"

"sin90\u00b0=1"

"I=\frac{\tau}{NAB}"

Put value

"I=\frac{18}{132\times(18.7\times10^{-2})^2\times0.242}"

"I=16.11A"